OpenStudy (anonymous):

A square has a length of 3 3/5 inches. What is the square's area and perimeter? WHich one is greater?

OpenStudy (anonymous):

square's area = length*length perimeter = 4*length

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

\[\left(3+\frac{3}{5}\right)^2=\frac{324}{25} \]\[4 \left(3+\frac{3}{5}\right)=\frac{72}{5}=\frac{72\ 5}{5\ 5}=\frac{360}{25} \]From Mathematica:\[\text{Reduce}\left[L^2\geq 4 L\right]\to L\leq 0\lor L\geq 4 \]The above says that for positive numbers, the area is greater than the perimeter when L, the square's side length, is greater than 4. When L = 4, the area and perimeter are equal to each other.